# Latent Diffusion Model-Enabled Real-Time Semantic Communication Considering Semantic Ambiguities and Channel Noises

### v1

@@ 1 Initialize $\bm{\omega}\leftarrow\bm{1}$ and $\bm{\nu}\leftarrow\bm{\gamma}^{L}$;
Return the parameters $\bm{\omega}$ of adaptor $g_{\bm{\omega}}(.)$
Algorithm 2 Online training algorithm of out-of-distribution adaptor $g_{\bm{\omega}}(.)$
In summary, when the SemCom system transmits data in domain, the parameters $\bm{\omega}$ of $g_{\bm{\omega}}(.)$ is equal to $\bm{1}$, while transmitting out-of-domain data, online learning Algorithm 2 will be activated. Due to the limited amount of training data and the fact that the given initial values of $\hat{\bm{z}}$ and $\bm{\omega}$ are close to the optimal values, this online-updated process is very fast. When the online-learning is completed, in order to not change the weights $\bm{\phi}^{\prime}$, $\bm{\psi}$, and $\bm{\theta}$ of robust encoder, generator, and LDM utilized for channel denoising by additional transfer learning, the adaptor is deployed at the receiver and defined as a dynamic lightweight nerual network. In other words, the parameter of the implemented adaptor $g_{\bm{\omega}}(.)$ can be dynamically changed through shared knowledge of SemCom system as shown in Fig. 1. Ultimately, the semantically consistent out-of-distribution data is reconstructed according to $\hat{\bm{z}}=g_{\bm{\omega}}(\bm{z})$.
@@ IV Latent Channel Denoising Diffusion Model
In this section, the wireless channel equalization under different conditions is firstly established in Subsection IV-A. The training objectives of original LDMs based on the received signals and its one-step real-time implementation are introduced in Subsection IV-B and IV-C.
@@ IV-A Wireless Channel Equalization
Through the joint encodeing of JSCC composed of robust encoder $E_{\bm{\phi}^{\prime}}(.)$ and generator $G_{\bm{\psi}}(.)$, the robust encoder encodes the original data $\bm{x}$ into a low-dimensional latent bottleneck $\bm{z}$. Assume that the latent vector $\bm{z}$ needs to reach the receiver through $k$ wireless channels, the transmitted complex latent signal is denoted by $\bm{z}_{c}\in\mathbb{C}^{k}$. At time $t$, the complex signal $\bm{y}_{c}$ received by the receiver on the $i$-th wireless channel can be represented as
$y_{c,i}=h_{c,i}z_{c,i}+n_{c,i},$ (15)
where $\bm{h}_{c}(t)=\sum_{i=1}^{k}\alpha_{i}e^{-j2\pi f\tau_{i}(t)}$, $\alpha_{i}$ is the signal amplitude of $i$-th path, $f$ is the carrier frequency, $\tau_{i}(t)$ denotes the phase shift, $\bm{n}_{c}\sim\mathcal{CN}(\bm{0},\sigma^{2}\bm{I})$ represents the complex Gaussian noises. Considering the effects of multipath fading and scattering, $h_{c,i}$ are independent and identically distributed (i.i.d.) Rician fading gains, which can be denoted by
$h_{c,i}=\sqrt{\frac{K}{K+1}}+\sqrt{\frac{1}{K+1}}h_{Rayleigh,i},$ (16)
where $h_{Rayleigh,i}$ are i.i.d. Rayleigh fading gains and $K$ is the ratio of direct radio waves’ power and non-direct radio waves’ power. When $K=\infty$, the wireless channels become additive white Gaussian noise (AWGN) channels, and the channels become Rayleigh channels when $K=0$. Optionally, minimum mean square error (MMSE) is utilized as a method for received signals equalization to avoid errors and improve efficiency. Consequently, the addressed signals can be donoted by
@@
For simplicity, the transmitted complex signals $\bm{z}_{c}$ can also be rewritten as $\bm{z}_{R}\in\mathbb{R}^{2k}$ in real-valued symbols, the equalization can be defined as $\bm{y}_{R}\in\mathbb{R}^{2k}$. In this way, the 1-st to $k$-th signals of $\bm{y}_{R}$ are
@@
where $\epsilon\sim\mathcal{N}(0,1)$. And the $k+1$-th to $2k$-th signals can defined as
@@
To this end, the diagonal matrices $\bm{H}_{z}$ and $\bm{H}_{n}$ can be defined as
@@
As a consequence, the conditional distribution of $\bm{y}_{R}$ under the estimated different wireless CSI, i.e., $\bm{h}_{c}$ and SNRs, is
$q_{\textrm{MMSE}}\left(\bm{y}_{R}|\bm{z}_{R},\bm{H}_{z},\bm{H}_{n}\right)=\mathcal{N}\left(\bm{y}_{R};\bm{H}_{z}\bm{z}_{R},\bm{H}^{2}_{n}\sigma^{2}\bm{I}\right).$ (22)
IV-B Latent Diffusion Model
The denoising task of receiver is to find the original transmitted signals $\bm{z}_{R}$ from transmitter given $\bm{y}_{R}$ and CSI. Accordingly, let $\bm{z}_{0}=\bm{H}_{z}\bm{z}_{R}$, the cross-entropy term in SemCom system model can be transformed from $\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x})}\left[-\log p_{\bm{\theta}}(\bm{z}|\bm{z}^{\prime})\right]$ to $\mathbb{E}_{q}\left[-\log p_{\bm{\theta}}(\bm{z}_{0}|\bm{y}_{R},\bm{H}_{z},\bm{H}_{n})\right]$. LDM is selected for wireless channel denoising as it has powerful capabilities to generate realistic data and much lower computational complexity than the original DMs. Let $\left\{\bm{z}_{t}\right\}^{t=T}_{t=0}$ be the noisy latent bottlenecks containing noises of different SNRs in the continuous time domain $t\in[0,T]$, where $\bm{z}_{0}$ is the starting latent vector. LDM defines a forward process through a unified stochastic differential equation (SDE)
@@ IV-B Latent Diffusion Model
where $\bm{u}(\bm{z},t)$ and $\bm{g}(t)$ are the drift and diffusion coefficients, and $\bm{w}_{t}$ is a standard Brownian motion. By considering the reverse process of SDE, the marginal distribution $p(\bm{z}_{t})$ follows the solution trajectory of the probability flow-ODE (PF-ODE)
@@
where $\nabla_{\bm{z}}\log p(\bm{z}_{t})$ denotes the score function. Accordingly, similar to EDM [58], considering the conditional distribution in Eq. (22), this paper sets $\bm{u}(\bm{z},t)=0$, $\bm{g}(t)=\sqrt{2t\bm{H}_{n}}$, and $\bm{\sigma}(t)=\bm{H}_{n}t$, where $t\in[0,T]$. When solving the revserse sampling trajectory, $t$ requires a discrete schedule $\left\{t_{n}\right\}^{n=N}_{n=0}$. Concretely, when $n=0$, $t_{0}=0$, and when $n\geq 1$, $t_{n}=\left(t_{1}^{1/\rho}+\frac{n-1}{N-1}\left(t_{N}^{1/\rho}-t_{1}^{1/\rho}\right)\right)^{\rho}$, where $\rho>0$. Moreover, unlike denoising diffusion probabilistic model (DDPM) [23], the utilized diffusion model adopts variance explosion (VE) strategy, and its associated forward process $\left\{\bm{z}_{t}\right\}^{t=T}_{t=0}$ can be written as
@@
In the reverse process, the denoising U-Net is usually utilized to predict $\bm{s}_{\bm{\theta}}(\bm{z},t)$ to approximate the score function $\nabla_{\bm{z}}\log p(\bm{z}_{t})$. Noise prediction model $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)$ are one of the most popular implementations of diffusion models, and $\bm{s}_{\bm{\theta}}(\bm{z},t)=-\frac{\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)}{\bm{H}_{n}t}$. As a consequence, the tranining objective of LDM is to minimize the loss
@@
$=\mathbb{E}_{\bm{z}_{R}\sim q_{\bm{\phi}^{\prime}}(\bm{\bm{z}_{R}|x}),\bm{\epsilon}_{1}\sim\mathcal{N}(\bm{0},\bm{I}),n\sim\mathcal{U}[1,N]}$
$\quad\left[\left\|\frac{\bm{\epsilon}_{\bm{\theta}}(\bm{H}_{z}\bm{z}_{R}+\bm{H}_{n}t_{n}\bm{\epsilon}_{1},t_{n})}{\bm{H}_{n}t_{n}}-\frac{\bm{\epsilon}}{\bm{H}_{n}t_{n}}\right\|^{2}_{2}\right]$
$\Leftrightarrow\mathbb{E}_{q}\left[\left\|\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)-\bm{\epsilon}\right\|^{2}_{2}\right].$
The PF-ODE defined in Eq. (24) can be rewritten as
@@
Similar to the channel denoising DM in [40], the variational evidence upper bound for wireless channel denoising tasks should consider different CSI conditions and practical DMs implementations, which is given by
$\mathcal{L}^{\prime}_{LDM}=\mathbb{E}_{q}\left[-\log p_{\bm{\theta}}\left(\bm{z}_{0}|\bm{y}_{R},\bm{H}_{z/n}\right)\right]$ (28)
$\leq\mathbb{E}_{q}\left[-\log\left(\frac{p_{\bm{\theta}}(\bm{z}_{t_{0}:t_{m}},\bm{y}_{R}|\bm{H}_{z/n})}{q(\bm{z}_{t_{1}:t_{m}},\bm{y}_{R}|\bm{z}_{0},\bm{H}_{z/n})}\right)\right]$
$=\mathbb{E}_{q}\bigg{[}\underbrace{D_{KL}\left(q_{\textrm{MMSE}}(\bm{y}_{R}|\bm{z}_{0},\bm{H}_{z/n})\parallel p(\bm{y}_{R}|\bm{H}_{z/n})\right)}_{\mathcal{L}_{\bm{y}_{R}}}$
$+\underbrace{D_{KL}\left(q(\bm{z}_{t_{m}}|\bm{y}_{R},\bm{z}_{0},\bm{H}_{z/n})\parallel p_{\bm{\theta}}(\bm{z}_{t_{m}}|\bm{y}_{R},\bm{H}_{z/n})\right)}_{\mathcal{L}_{t_{m}}}$
$+\sum^{m}_{n=2}\underbrace{D_{KL}\left(q(\bm{z}_{t_{n-1}}|\bm{z}_{t_{n}},\bm{z}_{0},\bm{H}_{z/n})\parallel p_{\bm{\theta}}(\bm{z}_{t_{n-1}}|\bm{z}_{t_{n}},\bm{H}_{z/n})\right)}_{\mathcal{L}_{t_{n-1}}}$
$-\underbrace{\log p_{\bm{\theta}}\left(\bm{z}_{t_{0}}|\bm{z}_{t_{1}},\bm{H}_{z/n}\right)}_{\mathcal{L}_{t_{0}}}\bigg{]},$
where $t_{m}$ is determined by $\mathop{\arg\min}_{t_{m}}\left|\sigma^{2}-t_{m}^{2}\right|$ and $m$ denotes the utilized denoising steps of LDM with respect to terms $\mathcal{L}_{\bm{y}_{R}}$ and $\mathcal{L}_{t_{m}}$. Evidently, assume that $N>m$, the terms $\mathcal{L}_{t_{n-1}}$ and $\mathcal{L}_{t_{0}}$ for wireless channel denoising are the sub-terms of the complete LDM training objectives. In other words, the selection of hyperparameters $T$ or $N$ should consider the possible worst SNRs of wireless channels. Ultimately, the transmitted latent vector $\bm{z}_{R}$ is given by $\bm{H}^{-1}_{z}\bm{z}_{0}$. Nonetheless, in the wireless communication scenarios with low SNRs, $m\gg 1$. As a result, LDM will execute $m$ times of noise predictions $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)$, i.e., the number of function evaluation (NFE) reachs $m$. Unfortunately, the varying fading wireless channels with uncertain SNRs bring significant uncertainties to the computational complexity of LDM, which undermines the possibility of implementing real-time wireless SemCom.
@@ IV-C End-to-End Consistency Distillation
The multi-step reverse sampling process of DMs brings the disadvantage of slow data generation speed. To overcome it, methods based on DDIM subsequence sampling [24], optimal reverse variances [59], LDMs [48, 60], SDE/ODE solvers [61, 62], and knowledge distillation [63, 46] have been proposed to optimize or accelerate the sampling process. In detail, the LDMs can significantly reduce the demensionality of input data, and some distillation based approaches only require a few steps or even one step to evaluate the output data without generation qulaity issues. Among these acceleration methods, the consistency model [46], as one of the distillation approaches, defines the consistency function $\bm{f}:(\bm{z}_{t},t)\mapsto\bm{z}_{\varepsilon}$ given a forward trajectory $\left\{\bm{z}_{t}\right\}_{t\in[\varepsilon,T]}$, where $\varepsilon=t_{1}\approx 0$. The consistency function assumes that for the input data on the same forward trajectory, the output of the neural network parameterized function points to the same generated data, which is given by
@@
where $\hat{\bm{\theta}}$ is the nerual network parameters of consistency model. By observing function $\bm{F}_{\hat{\bm{\theta}}}(\bm{z}_{t},t)$, it can be implemented by directly training a neural network to map noisy data $\left\{\bm{z}_{t}\right\}_{t\in(\varepsilon,T]}$ to $\bm{z}_{\varepsilon}$. Nevertheless, this type of network is similar to the denoising autoencoder, which has poor performances on wireless channel denoising task. Accordingly, the consistency function $\bm{f}_{\hat{\bm{\theta}}}(\bm{z}_{t},t)$ can be obtained by distilling the pretrained original LDM $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)$.
Figure 4: In the proposed latent generative model, data is mapped into latent space via robust encoder $q_{\bm{\phi}^{\prime}}(\bm{z}_{0}|\bm{x})$ and the forward diffusion model is applied in latent space ($\bm{z}_{0}\rightarrow\bm{z}_{T}$). Then, EECD maps noisy received signals to denoised latent vector ($\bm{z}_{t_{m}}\rightarrow\bm{z}_{\varepsilon}$) and decoder will generate data with desired semantic meaning by $p_{\bm{\psi}}(\bm{x}|\bm{z}_{\varepsilon})$.
which is also called denoising diffusion implicit model (DDIM) [24] and every step’s NFE equals to 1. However, the actual values of the difference $\frac{\bm{z}_{t_{n+1}}-\bm{z}_{t_{n}}}{t_{n+1}-t_{n}}$ are closer to the derivatives between $\bm{z}_{t_{n+1}}$ and $\bm{z}_{t_{n}}$, rather than the dervatives at $\bm{z}_{t_{n+1}}$. To this end, the Heun solver in EDM is adopted [58], which is denoted by
where the $\bm{z}_{t_{n}}$ on the right side of Eq. (31) can be approximated by $\tilde{\bm{z}}^{\bm{\theta}}_{t_{n}}$ in Eq. (30). Consequently, the estimation of $\bm{z}_{t_{n}}$ is given by
where the NFE equals to 2. According to the definition of the consistency function, the function $\bm{f}_{\hat{\bm{\theta}}}(\bm{z}_{t},t)$ should has the same output for adjacent data points $(\bm{z}_{t_{n+1}},t_{n+1})$ and $(\hat{\bm{z}}^{\bm{\theta}}_{t_{n}},t_{n})$ on the same reverse trajectory, i.e., the loss of the consistency model is
where $\hat{\bm{\theta}}^{-}$ denotes the running average of the past values of $\hat{\bm{\theta}}$ during optimization. Nonetheless, the goal of wireless SemCom is to accurately reconstruct the transmitted data in real-time manner in the receiver side. Inspired by that, the CD loss in Eq. (33) can be changed to the loss of EECD, which is given by
where $\bm{d}(\cdot,\cdot)$ is denoted by $\mathcal{L}_{2}$ distance in non-images datasets, and learned perceptual image path similarity distance (LPIPS) $\mathcal{L}_{LPIPS}$ [64] in image datasets.

### v4

@@ 1 Initialize $\bm{\omega}\leftarrow\bm{1}$ and $\bm{\nu}\leftarrow\bm{\gamma}^{L}$;
Return the parameters $\bm{\omega}$ of adaptor $g_{\bm{\omega}}(\cdot)$
Algorithm 2 Online training algorithm of out-of-domain adaptor $g_{\bm{\omega}}(\cdot)$
In summary, when the SemCom system transmits data in domain, the parameters $\bm{\omega}$ of $g_{\bm{\omega}}(\cdot)$ is equal to $\bm{1}$, while transmitting data with significant reconstruction errors inferred by the decoder deployed at the transmitter, online learning Algorithm 2 will be activated to improve the decoded data’s quality. Due to the limited amount of training data and the fact that the given initial values of $\hat{\bm{z}}$ and $\bm{\omega}$ are close to the optimal values, this online-updated process will be very fast. When the online-learning is completed, in order to not change the weights $\bm{\phi}^{\prime}$, $\bm{\psi}$, and $\bm{\theta}$ of robust encoder, generator, and LDM utilized for channel denoising, the adaptor is deployed at the receiver and defined as a dynamic lightweight nerual network. In other words, the parameters of the implemented adaptor $g_{\bm{\omega}}(\cdot)$ can be dynamically changed in the proposed SemCom system as shown in Fig. 1. Ultimately, the semantically consistent out-of-domain data is reconstructed according to $\hat{\bm{z}}=g_{\bm{\omega}}(\bm{z})$.
@@ IV Latent Channel Denoising Diffusion Model
In this section, the wireless channel equalization under different conditions is firstly established in Subsection IV-A. The training objectives of original LDMs based on the received signals and its one-step real-time implementation are introduced in Subsections IV-B and IV-C.
@@ IV-A Wireless Channel Equalization
Optionally, minimum mean square error (MMSE) [57] is usually utilized as a method for received signals equalization to avoid errors and improve efficiency. Consequently, let $\bm{h}_{c}=[h_{c,1},\cdots,h_{c,k}]$ and $\bm{n}_{c}=[n_{c,i},\cdots,n_{c,k}]$ be the channel state and noises, the addressed signals for the received signals $\bm{z}^{\prime}=\bm{y}_{c}$ defined in Section II can be denoted by
@@
For simplicity, the transmitted complex signals $\bm{z}_{c}$ can also be rewritten as $\bm{z}_{R}\in\mathbb{R}^{2k}$ in real-valued symbols, the output of equalization can be also decoupled as corresponding real-valued $\bm{y}_{R}\in\mathbb{R}^{2k}$. In this way, the 1-st to $k$-th components of $\bm{y}_{R}$ are
@@
where $\epsilon\sim\mathcal{N}(0,1)$. And the $k+1$-th to $2k$-th components can defined as
@@
To this end, the known diagonal CSI matrix $\bm{H}_{z}$ and noise coefficient matrix $\bm{H}_{n}$ can be defined as
@@
As a consequence, the conditional distribution of $\bm{y}_{R}$ under the estimated wireless CSI, i.e., $\bm{h}_{c}$ and SNRs, is
$q_{\textrm{MMSE}}\left(\bm{y}_{R}|\bm{z}_{R},\bm{H}_{z},\bm{H}_{n}\right)=\mathcal{N}\left(\bm{y}_{R};\bm{H}_{z}\bm{z}_{R},\bm{H}^{2}_{n}\sigma^{2}\bm{I}\right),$ (19)
which means the received signals is affected by the channel’s fading gains and noises. Especially, $\bm{H}_{z}=\bm{H}_{n}=\bm{I}\in\mathbb{R}^{2k\times 2k}$ under AWGN channel.
IV-B Multi-Step Latent Diffusion Model
The denoising task of receiver is to find the original transmitted signals $\bm{z}_{R}$ from transmitter given $\bm{y}_{R}$ and CSI. Accordingly, let $\bm{z}_{0}=\bm{H}_{z}\bm{z}_{R}$, the cross-entropy term in SemCom system model can be transformed from $\mathbb{E}_{q}\left[-\log p_{\bm{\theta}}(\bm{z}|\bm{z}^{\prime})\right]$ to $\mathbb{E}_{q}\left[-\log p_{\bm{\theta}}(\bm{z}_{0}|\bm{y}_{R},\bm{H}_{z},\bm{H}_{n})\right]$. LDM is selected for wireless channel denoising as it has powerful capabilities to generate realistic data and much lower computational complexity than the original DMs. Let $\left\{\bm{z}_{t}\right\}^{t=T}_{t=0}$ be the noisy latent bottlenecks containing noises of different SNRs in the continuous time domain $t\in[0,T]$, where $\bm{z}_{0}$ is the starting latent vector. LDM defines a forward process through a unified stochastic differential equation (SDE)
@@ IV-B Multi-Step Latent Diffusion Model
where $\bm{u}(\bm{z},t)$ and $\bm{g}(t)$ are the drift and diffusion coefficients, and $\bm{w}_{t}$ is a standard Brownian motion. By considering the reverse process of SDE, the marginal distribution $p(\bm{z}_{t})$ follows the solution trajectory of the probability flow-ordinary differential equation (PF-ODE)
@@
where $\nabla_{\bm{z}}\log p(\bm{z}_{t})$ denotes the score function. Accordingly, similar to Elucidated DM (EDM) [58], considering the conditional distribution in Eq. (19), this paper sets $\bm{u}(\bm{z},t)=0$, $\bm{g}(t)=\sqrt{2t\bm{H}_{n}}$, and $\bm{\sigma}(t)=\bm{H}_{n}t$, where $t\in[0,T]$. When solving the reverse sampling trajectory, $t$ requires a discrete schedule $\left\{t_{n}\right\}^{n=N}_{n=0}$. Concretely, when $n=0$, $t_{0}=0$, and when $n\geq 1$, $t_{n}=\left(t_{1}^{1/\rho}+\frac{n-1}{N-1}\left(t_{N}^{1/\rho}-t_{1}^{1/\rho}\right)\right)^{\rho}$, where $\rho>0$. Moreover, unlike denoising diffusion probabilistic model (DDPM) [21], the utilized diffusion model adopts variance explosion (VE) strategy, and its associated forward process $\left\{\bm{z}_{t}\right\}^{t=T}_{t=0}$ can be written as
@@
In the reverse process, the denoising U-Net is usually utilized to predict approximation function $\bm{s}_{\bm{\theta}}(\bm{z},t)$ to approximate the score function $\nabla_{\bm{z}}\log p(\bm{z}_{t})$. Noise prediction model $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)$ is one of the most popular implementations of diffusion models, and $\bm{s}_{\bm{\theta}}(\bm{z},t)=-\frac{\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)}{\bm{H}_{n}t}$ [59]. As a consequence, the tranining objective of LDM is to minimize the distance between noise prediction $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)$ and actual noise $\bm{\epsilon}$ [59]
@@
$=\mathbb{E}_{\bm{z}_{R},\bm{\epsilon}_{1},n}\left[\left\|\frac{\bm{\epsilon}_{\bm{\theta}}(\bm{H}_{z}\bm{z}_{R}+\bm{H}_{n}t_{n}\bm{\epsilon}_{1},t_{n})}{\bm{H}_{n}t_{n}}-\frac{\bm{\epsilon}}{\bm{H}_{n}t_{n}}\right\|^{2}_{2}\right]$
$\Leftrightarrow\mathbb{E}_{q}\left[\left\|\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)-\bm{\epsilon}\right\|^{2}_{2}\right],$
where $\bm{\epsilon}_{1}\sim\mathcal{N}(\bm{0},\bm{I})$ and $n\sim\mathcal{U}[1,N]$. Considering aforementioned conditions and settings, the PF-ODE defined in Eq. (21) can be rewritten as
@@
Similar to the channel denoising DM in [36], wireless channel denoising task is a subprocess of whole diffusion reverse process. The denoising start point $t_{m}$ should be determined by $\mathop{\arg\min}_{t_{m}}\left|\sigma^{2}-t_{m}^{2}\right|$ with known $\sigma^{2}$ and $m$ denotes the utilized denoising steps of pretrained LDM. Consequently, The selection of hyperparameters $N$ and $T$ should consider the worst-case SNRs to make the channel denoising objective a sub-term of the DM training objective. Ultimately, the transmitted latent vector $\bm{z}_{R}$ is given by $\bm{H}^{-1}_{z}\bm{z}_{0}$. Nonetheless, in the wireless communication scenarios with large noise variance $\sigma^{2}$, $m\gg 1$ according to the designed discrete schedule $\left\{t_{n}\right\}^{n=N}_{n=0}$. As a result, LDM will execute $m$ times of noise predictions $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)$, i.e., the number of function evaluations (NFE) will reach $m$. Unfortunately, the varying fading wireless channels with uncertain SNRs bring significant uncertainties to the computational complexity of LDM, which undermines the possibility of implementing real-time SemCom.
@@ IV-C End-to-End Consistency Distillation
The multi-step reverse sampling process of DMs brings the disadvantage of slow data generation speed. To overcome it, methods based on denoising diffusion implicit model (DDIM) subsequence sampling [22], optimal reverse variances [38], LDMs [40], SDE/ODE solvers [59], and knowledge distillation [42] have been proposed to optimize or accelerate the sampling process. In detail, the LDMs can significantly reduce the demensionality of input data, and some distillation based approaches only require a few steps or even one step to evaluate the output data without generation quality issues. Among these acceleration methods, the consistency model [42], as one of the distillation approaches, defines the consistency function $\bm{f}:(\bm{z}_{t},t)\mapsto\bm{z}_{\varepsilon}$ given a forward trajectory $\left\{\bm{z}_{t}\right\}_{t\in[\varepsilon,T]}$, where $\varepsilon=t_{1}\approx 0$. The consistency function assumes that for the input data on the same forward trajectory, the output of the neural network parameterized function points to the same generated data, which is given by
@@
where $\hat{\bm{\theta}}$ is the neural network parameters of consistency model. By observing function $\bm{F}_{\hat{\bm{\theta}}}(\bm{z}_{t},t)$, it can be implemented by directly training a neural network to map noisy data $\left\{\bm{z}_{t}\right\}_{t\in(\varepsilon,T]}$ to $\bm{z}_{\varepsilon}$. Accordingly, the consistency function $\bm{f}_{\hat{\bm{\theta}}}(\bm{z}_{t},t)$ can be obtained based on EDM architecture by distilling the pretrained original LDM $\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t)$.
Figure 4: In the proposed SemCom model, data is mapped into latent space via robust encoder $q_{\bm{\phi}^{\prime}}(\bm{z}_{0}|\bm{x})$. Then, EECD maps noisy received signals to denoised latent vector ($\bm{z}_{t_{m}}\rightarrow\bm{z}_{\varepsilon}$) and decoder will generate data with desired semantic meaning by $p_{\bm{\psi}}(\bm{x}|\bm{z}_{\varepsilon})$.